# Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation

### v1

@@ 3.4 Knowledge-Based Dialogue Response Generation
Prompt-based generation can effectively improve factuality with training, while it is unable to capture the connectivity information within a Knowledge Graph. To address this limitation, we encode the knowledge graph with GNN (Graph Netural Network). We first convert the sorted triples $f^{{}^{\prime}}_{sorted}$ into a graph $\mathcal{G}$. Then the graph and dialogue encoding are defined as follows:
$v=E_{graph}(\mathcal{G})$ (8)
$z=\mathcal{M}(c^{\prime},x^{\prime})$ (9)
where $E_{graph}$ is GNN. $\mathcal{M}$ represents an arbitrary LLM. $v$ and $z$ denote the graph and text embedding, respectively.
We generate the dialogue response by decoding the concatenation of the graph and text embedding, formulated as follows:
$s_{g}=Decoder([v;z])$ (10)
where ; denotes concatenation and $s_{g}$ is the response generated by graph-based method. To preserve the capability of the LLM, we combine LoRA with cross-entropy to train our graph model.
3.5 Evaluation Metrics on Fact
Fact score Min et al. (2023) measures the fact consistency for long-form text. As we mentioned in Section 2.2, it does not consider the dialogue history and the possibility that the generated response may lack the knowledge source to support it.
The atomic facts and fact verification prompts of fact score are modified in this paper. In the revised atomic fact-splitting prompt, only complete sentences are permitted to be split, and the model is prohibited from adding any additional information to the atomic facts. We incorporate dialogue history and introduce the “Not Enough Information” label for the revised fact verification prompt. If an atomic fact is not a factual claim or lacks direct support from any known sources, the model will output ‘Not Enough Information”. The detailed prompts for atomic fact-splitting and fact verification are provided in Table 8 of Appendix A.
@@ 3.5 Evaluation Metrics on Fact
Since the original fact score framework does not specify how to get knowledge sources for dialogue, we extract entities from atomic facts and query as Wikipedia titles to search related passages as knowledge sources.
@@ 4.1 Dataset
Two public dialogue datasets, OpendialKG Moon et al. (2019) and HybriDialogue Nakamura et al. (2022), are used to evaluate our methods. The two datasets both provide response references but not factual ones. See Appendix B for the details and examples of datasets.
@@ 4.2 Conventional Evaluation Metrics
We aim to evaluate factuality and quality for dialogue response generation. BLEU Reiter (2018), ROUGE-L Lin (2004), Perplexity Jelinek et al. (1977) and F1 score Nan et al. (2021) are employed to evaluate the quality and fact consistency of the generated responses. For more details, please refer to Appendix E.
@@ 4.3 Modified Evaluation Metrics
We adopt both the adapted fact score and the NEIP (Not Enough Info Proportion) to comprehensively evaluate factuality. We also manually assess fact score with humans, as shown in Table 1. See the Appendix F for more annotation details.
Datasets Agreement Cohen’s Kappa
@@
Table 1: The agreement and Cohen’s Kappa between the ground truth and LLM outputs regarding the fact score.
@@
ChatGLM-6B 1.42 14.24 12.98 70.91 50.13 15.14
   +BM25 1.38 14.02 13.15 70.99 49.07 14.92
   +KAPING 1.31 13.93 13.15 72.37 48.74 15.34
   +Triton((Ours) 1.37 14.07 13.31 74.64 48.73 15.40
   +Triton-C(Ours) 1.34 13.47 14.41 76.46 37.93 15.33
[…]
   +BM25 2.65 12.90 18.92 69.32 56.09 16.25
   +KAPING 2.56 12.72 18.18 68.34 55.79 16.03
   +Triton(Ours) 2.60 12.66 18.89 71.56 54.28 16.09
   +Triton-C(Ours) 2.19 11.27 20.22 74.25 28.53 15.00
[…]
   +BM25 2.68 12.71 23.18 72.06 31.98 13.69
   +KAPING 2.91 12.82 23.22 73.00 32.07 13.85
   +Triton(Ours) 3.16 13.06 23.07 76.20 31.42 14.50
   +Triton-C(Ours) 2.33 10.69 20.03 77.77 20.35 13.78
[…]
Triton-X(Ours) 3.99 25.86 14.43 87.73 59.20 13.75
Table 2: The experimental result with various methods on the OpendialKG dataset. The column heads with * indicate our primary metrics for evaluating factual accuracy. Ours means generating responses with our proposed methods. The bold number is the best result within each model. PPL indicates Perplexity and a lower PPL means more fluency. NEIP denotes the “not enough information” proportion, representing the percentage of atomic facts that either lack direct knowledge support or fail to qualify as factual claims.
@@ 4.4 Baseline Methods
Several widely recognised LLMs, such as ChatGLM Zeng et al. (2022) and Flan-T5 Chung et al. (2022), have been selected for this work. Furthermore, we compare our proposed Triton, Triton-C and Triton-X with BM25 Robertson et al. (2009), KAPING Baek et al. (2023), G-Retriever He et al. (2024), and a basic dialogue generation model. Additional details of LLMs can be found in the Appendix G.
[…]
ChatGLM-6B 4.62 24.03 23.66 46.58 58.19 39.57
   +BM25 4.68 24.19 23.77 53.32 55.18 40.26
   +KAPING 4.56 24.07 23.58 55.84 54.51 40.99
   +Triton(Ours) 4.71 24.43 23.95 58.15 53.07 40.96
   +Triton-C(Ours) 4.53 24.23 26.03 59.10 50.99 40.92
[…]
   +BM25 8.71 25.81 32.79 67.42 47.43 38.50
   +KAPING 8.92 26.43 31.96 68.92 46.56 40.41
   +Triton(Ours) 8.90 26.14 30.81 70.79 43.52 39.70
   +Triton-C(Ours) 8.20 24.63 34.01 74.41 33.68 40.63
[…]
   +BM25 11.08 29.75 38.89 60.40 48.79 44.65
   +KAPING 10.95 29.44 37.88 63.84 48.94 44.13
   +Triton(Ours) 10.93 29.30 41.35 66.62 46.49 43.91
   +Triton-C(Ours) 9.89 28.19 32.48 69.34 39.23 43.65
[…]
Triton-X(Ours) 20.85 40.77 39.12 73.00 54.87 43.98
Table 3: The experimental result on the HybriDialogue dataset with different baselines. The best results within each model are displayed in bold.
@@ 4.5 Experimental Setup
We illustrate the details of the experimental setup in Appendix D.
@@ 4.6 Experimental Result and Analysis
Tables 2 and 3 show the experimental results of different methods on two public datasets. We analyze the generated responses from the perspectives of text quality and factual consistency.
@@
With the KG, the BLEU-4 and ROUGE-L scores fluctuate across two datasets. Our proposed Triton, the knowledge retriever, generally perform better than other knowledge-enhanced baselines, while we see a drop with Triton-C, the combination of dialogue rewrite and knowledge retriever. Our proposed framework, Triton-X, which consists of three components, achieves the highest BLEU and ROUGE-L scores, surpassing the SOTA G-retriever and demonstrating the superiority of our method.
There is a slight growth in PPL of smaller language models like ChatGLM and Flan-T5-Large: incorporating the KG decreases text fluency. After analysis, we found that the LLMs sometimes can not understand the KG well and thus generate unpolished responses. We also utilise Case 1 from Section 4.9 to illustrate the factors contributing to low PPL.
Overall, knowledge-augmented methods show variable performance across different LLMs in terms of BLEU-4 and ROUGE-L scores, and they tend to increase PPL slightly for smaller language models. Our proposed Triton-X gains the best performance in BLEU-4 on both datasets and ROUGE-L on HybriDialogue, indicating a higher quality in generating responses.
Factuality
With retrieved methods, the fact score rises in all LLMs, and Triton has a higher increase than KAPING and BM25. From Triton-C, we can see a continuous improvement in fact score, which means the rewritten dialogue also contributes a slight improvement to the fact score.
The NEIP is the proportion of atomic facts that can not be directly found in external knowledge to support or which are not a factual claim, and the lower one indicates the responses contain more verifiable facts. Compared to baseline LLMs, knowledge-augmented methods generally help reduce this proportion and produce more verifiable facts. From the tables, we observe a consistent decrease in NEIP for the Triton compared to KAPING and BM25. Furthermore, we noticed a significant improvement in NEIP with the rewritten dialogue. This suggests that rewriting the dialogue can substantially enhance factuality.
It is worth noting that the F1 score is not always consistent with the fact score. For instance, the trend of the F1 score shown in Flan-T5-XXL is almost opposite to the fact score on the OpendialKG dataset. One reason is that the reference responses lack explicit answers. Another reason is that the F1 score ignores the semantic aspects, and two sentences with the same entities can express contrary meanings.
In short, knowledge-augmented methods improve the factuality of LLMs, including the fact score and NEIP. Our proposed Triton series methods perform better than G-Retriever, KAPING and BM25 in the above metrics.
@@
We conduct experiments with different LLMs. The Flan-T5 series LLMs (Large and XXL) are beneficial for us in analysing the correlation between size and performance.
As the model scale grows, fact score and NEIP exhibit differing trends across the two datasets—factuality declines on HybriDialogue, whereas the OpendialKG dataset shows the opposite pattern. The factuality is not always consistent with the model size, which contradicts our intuition.
HybriDialogue OpendialKG
[…]
w/o Retriever -1.1 -0.11 -1.1 +1.2
w/o Rewrite -2.3 +0.48 -1.46 +0.44
w/o Graph -3.41 -0.03 -1.00 -1.52
Table 4: Ablation study of Triton-X, and w/o means removing corresponding components.
@@ 4.7 Ablation Study
We conduct an ablation study with our proposed framework, shown in Table 4. We remove our components one by one from Triton-X to observe the performance. We only report factually related metrics as they are our aim. The decline in performance after removing either the rewrite or the retriever component highlights their importance.
@@ 4.8 Human Evaluation Results
ChatGLM-6B 1.90 1.99 1.71 1.99 1.91 1.11
+bm25 1.85 1.97 1.75 1.99 1.89 1.09
+SenEmb 1.84 1.98 1.71 1.98 1.95 1.18
+Triton 1.93 1.98 1.76 2.00 1.94 1.11
+Triton-C 1.94 1.96 1.85 1.98 1.97 1.28
Flan-T5-XXL 1.62 1.74 1.74 1.56 2.00 1.36
+bm25 1.67 1.72 1.80 1.38 1.94 1.49
+SenEmb 1.59 1.80 1.77 1.46 1.99 1.53
+Triton 1.59 1.81 1.77 1.52 1.98 1.39
+Triton-C 1.51 1.83 1.76 1.47 1.97 1.65
G-retriever 1.92 1.82 1.84 1.96 1.96 0.71
TritonX 1.96 1.81 1.82 2.00 2.00 0.60
@@
We report the human evaluation with the perspective: coherence, fluency and informativeness in Table 5. Each aspect is rated as one of three levels: positive, neutral, or negative (e.g., coherent, neutral, incoherent).
Our proposed Triton series methods demonstrate notable improvements in coherence across various settings, with the exception of the Flan-T5-XXL model. In terms of fluency, our methods achieve the highest scores for Flan-T5-XXL on HybriDialogue, as well as for ChatGLM and graph-based methods on OpendialKG. Overall, the proposed methods maintain strong performance in fluency and coherence. Additionally, the Triton series generally enhances informativeness.

### v2

@@ 3.3 Graph Knowledge-Driven Response Generation
Textualised graph-based generation can effectively improve factuality without training, while it is unable to capture the connectivity information within a KG and furthermore affects the factuality of the response. To address this limitation, we encode the knowledge graph based on GNNs (Graph Neural Networks) (Kipf 2016).
To generate responses with a graph, we aim to encode the dialogue sense graph $\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}$ into a graph-level representation. We firstly encode its nodes $\mathcal{V}^{{}^{\prime}}$:
$\mathbf{H}^{0}=\text{GraphEmb}(\mathcal{V}^{\prime})\in\mathbb{R}^{|\mathcal{V}^{\prime}|\times d},$ (5)
where $d$ is the embedding dimension. The matrix $\mathbf{H}^{0}$ consists of node embeddings produced by the Graph Embedding Layer (GraphEmb). To get the higher node representation, we utilise a layer based on multi-head attention (Vaswani et al. 2017), which is formulated as follows:
$\mathbf{H}^{l+1}=\text{GTC-LN}(\mathbf{H}^{l})\in\mathbb{R}^{|\mathcal{V}^{\prime}|\times d},$ (6)
where we employ Graph TransformerConv (Shi et al. 2020) with Layer Norm (GTC-LN) to encode $\mathbf{H}^{0}$ to higher node representations $\mathbf{H}^{l+1}$, which can capture more global structural information and lay the foundation for generating high factually correct responses.
$\mathbf{h}_{\text{mean}}=\frac{1}{|\mathcal{V}^{\prime}|}\sum_{v\in\mathcal{V}^{\prime}}\mathbf{H}^{l+1}_{v}\in\mathbb{R}^{d},$ (7)
$\mathbf{h}_{\text{graph}}=\text{GKProj}(\mathbf{h}_{\text{mean}})\in\mathbb{R}^{d}.$ (8)
Here we mean pooling the node embeddings to get a global-level representation $\mathbf{h}_{\text{mean}}$, and further, a Graph Knowledge Projector (GKProj), implemented by a feed-forward neural network, is used to transfer the $\mathbf{h}_{\text{mean}}$ to a graph-level representation $\mathbf{h}_{\text{graph}}$. The projector serves as a bridge between the graph and response generation.
$\mathbf{z}=\text{TextEmb}(\mathcal{U})\in\mathbb{R}^{d}.$ (9)
Here we encode the utterances $\mathcal{U}$ to a text embedding $\mathbf{z}$ by the Text Embedding Layer (TextEmb) of an LLM.
We generate the token $\text{U}^{*}_{ti}$ by sampling the probability of $\text{U}^{*}_{ti}$ conditioned on generated tokens $\text{U}^{*}_{t,<i}$, $\mathbf{h}_{\text{graph}}$, $\mathbf{z}$ and the parameter $\theta$, formulated as follows:
$\text{U}^{*}_{ti}\sim P(\text{U}^{*}_{ti}\mid\text{U}^{*}_{t,<i},\mathbf{h}_{\text{graph}},\mathbf{z};\theta).$ (10)
We train the model using the following cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{T}\sum_{c=1}^{C}y_{i}^{(c)}\log p_{i}^{(c)},$ (11)
where $p_{i}^{(c)}=P(\text{U}^{*}_{ti}=c\mid\text{U}_{t,<i},\mathbf{h}_{\text{graph}},\mathbf{z};\theta)$ is the model-predicted probability over vocabulary class $c$ at step $i$. To retain the pretrained knowledge of the LLM, we adopt LoRA combined with cross-entropy to train our graph model. We refer the generation $\text{U}^{*}_{t}$ to GA-DRG.
3.4 Dialogue Fact Score
We propose the Dialogue Fact Score that incorporates dialogue context and explicitly handles cases where supporting information is unavailable.
We redesign the atomic fact extraction and verification prompts used in computing FactScore. In the redesigned atomic fact extraction prompt, only complete sentences are eligible for splitting, and the model is instructed not to introduce any additional information when generating.
For fact verification, we revise the prompt to incorporate dialogue history and introduce a new class: “Not Enough Information”. If an atomic fact is not a verifiable factual claim or lacks direct support from any known source, the model outputs “Not Enough Information.”
@@ 3.4 Dialogue Fact Score
We leverage the LLaMA 3.3 70B Instruct model to compute the dialogue fact score. Since the FactScore does not define the knowledge retrieval process for dialogue, we extract entities from atomic facts and dialogue as Wikipedia titles to retrieve passages as supporting knowledge.
@@ 4.1 Dataset
Two public knowledge-driven and English dialogue datasets, OpendialKG (Moon et al. 2019) and HybriDialogue (Nakamura et al. 2022), are used in our work. OpenDialKG consists of a recommendation task related to movies and books and a chit-chat task related to sports and music. HybriDialogue is an open-domain and information-seeking dialogue dataset.
@@ 4.2 Conventional Evaluation Metrics
We aim to evaluate factuality and quality for dialogue response generation.
BLEU (Reiter 2018), ROUGE-L (Lin 2004), Perplexity (Jelinek et al. 1977) and F1 score (Nan et al. 2021) are employed to evaluate the quality and fact consistency of the generated responses automatically.
For human evaluation, Coherence checks if the response makes sense and matches the previous context. Fluency looks at whether the response is grammatically correct and sounds natural. Informativeness measures if the response adds useful or new information to the conversation. We randomly extracted 50 samples for each model and recruited two annotators for annotation. HybriDialogue achieves Agreement/Kappa scores of 0.83/0.41 for Coherence, 0.90/0.53 for Fluency, and 0.89/0.53 for Informativeness, while OpendialKG scores 0.83/0.47, 0.96/0.27, and 0.80/0.60 on the same metrics, respectively. However, Kappa is less effective in reflecting the true agreement under imbalanced conditions. In the fluency task of OpendialKG, one class accounts for approximately 97% of the data, making it hard for kappa to accurately measure the agreement.
@@ 4.3 Modified Evaluation Metrics
We adopt both the dialogue fact score and the NEIP (Not Enough Info Proportion) to evaluate factuality comprehensively. Two human annotators manually assessed the dialogue fact score with 100 random samples on two datasets, respectively, as shown in Table 1. The Cohen’s Kappa scores are both above 0.6, indicating substantial agreement between humans and the dialogue score.
Datasets Raw Agreement Cohen’s Kappa
@@
Table 1: Raw agreement and Cohen’s Kappa between ground truth and LLM dialogue fact scores.
OpendialKG HybriDialogue
@@
ChatGLM-6B 70.91 50.13 15.14 12.98 1.42 14.24 46.58 58.19 39.57 23.66 4.62 24.03
+BM25 70.99 49.07 14.92 13.15 1.38 14.02 53.32 55.18 40.26 23.77 4.68 24.19
+KAPING 72.37 48.74 15.34 13.15 1.31 13.93 55.84 54.51 40.99 23.58 4.56 24.07
+TG-DRG-NR 74.64 48.73 15.40 13.31 1.37 14.07 58.15 53.07 40.96 23.95 4.71 24.43
+TG-DRG-R 76.46 37.93 15.33 14.41 1.34 13.47 59.10 50.99 40.92 26.03 4.53 24.23
Flan-T5-Large 67.76 54.94 16.35 18.37 2.57 13.00 60.04 50.48 40.04 30.89 9.61 27.12
+BM25 69.32 56.09 16.25 18.92 2.65 12.90 67.42 47.43 38.50 32.79 8.71 25.81
+KAPING 68.34 55.79 16.03 18.18 2.56 12.72 68.92 46.56 40.41 31.96 8.92 26.43
+TG-DRG-NR 71.56 54.28 16.09 18.89 2.60 12.66 70.79 43.52 39.70 30.81 8.90 26.14
+TG-DRG-R 74.25 28.53 15.00 20.22 2.19 11.27 74.41 33.68 40.63 34.01 8.20 24.63
Flan-T5-XXL 69.29 36.62 14.07 23.37 3.44 13.63 53.06 50.62 44.07 37.47 11.34 29.62
+BM25 72.06 31.98 13.69 23.18 2.68 12.71 60.40 48.79 44.65 38.89 11.08 29.75
+KAPING 73.00 32.07 13.85 23.22 2.91 12.82 63.84 48.94 44.13 37.88 10.95 29.44
+TG-DRG-NR 76.20 31.42 14.50 23.07 3.16 13.06 66.62 46.49 43.91 41.35 10.93 29.30
+TG-DRG-R 77.77 20.35 13.78 20.03 2.33 10.69 69.34 39.23 43.65 32.48 9.89 28.19
G-Retriever 84.79 62.65 11.70 14.98 3.79 26.00 70.29 55.77 42.05 49.66 20.36 39.53
GA-DRG 87.73 59.20 13.75 14.43 3.99 25.86 73.00 54.87 43.98 39.12 20.85 40.77
Table 2: The experimental results on the OpendialKG and HybriDialogue dataset. The column heads with * indicate our primary metrics for evaluating factuality. Bold models mean generating responses with our proposed methods. The bold number is the best result within each category. NEIP denotes the “not enough information” proportion, representing the percentage of atomic facts that either lack direct knowledge support or fail to qualify as factual claims. PPL indicates Perplexity, and a lower PPL means more fluency. Fact, BLEU and ROUGE denote Dialogue Fact Score, BLEU-4 and ROUGE-L, respectively.
@@ 4.4 Baseline Methods
Several widely recognised LLMs, such as ChatGLM (Zeng et al. 2022) and Flan-T5 (Chung et al. 2022), have been selected for this work. Furthermore, we compare our proposed TG-DRG-NR, TG-DRG-R and GA-DRG with BM25 (Robertson, Zaragoza et al. 2009), KAPING (Baek, Aji, and Saffari 2023), G-Retriever (He et al. 2024).
@@ 4.5 Experimental Setup
Except for the GNN settings where the rank and alpha parameters are set to 8 and 16, respectively, the rank and alpha parameters for LoRA (Hu et al. 2021) are consistently set to 32. As G-retriever is implemented using Llama2 7B (Touvron et al. 2023) with LoRA, we adopt the same model under identical settings to ensure fairness. We utilise the multi-qa-mpnet-base-dot-v1 version for the baseline KAPING. All experiments were run a single time.
We utilised the evaluate package to compute BLEU-4 and ROUGE-L scores. Besides, the PPL is calculated using GPT-2 (Radford et al. 2019).
We used an A100 GPU for fine-tuning and inference. We utilised two A100 GPUs for the computation of the dialogue fact score. All A100 GPUs we used in this work have 80GB of memory.
@@ 4.6 Experimental Result and Analysis
We compare our proposed TG-DRG, which leverages a textualised graph, and GA-DRG, which utilises a GNN-based graph encoding approach, against LLM-based baselines, as well as BM25, KAPING, and G-Retriever. Table 2 shows the experimental results of different methods on two public datasets. We analyze the generated responses from the perspectives of factuality, text quality and model scale.
Factuality
With knowledge augmented methods, the dialogue fact score rises in all LLMs, and our TG-DRG-NR has a higher increase than KAPING and BM25. From TG-DRG-R, we can see a continuous improvement in dialogue fact score, which means the dialogue reformulation also contributes a slight improvement to the factuality.
The NEIP is the proportion of atomic facts that can not be directly found in external knowledge to support or which are not factual claims, and the lower one indicates the responses contain more verifiable facts. Compared to baseline LLMs, knowledge-augmented methods generally help reduce this proportion and produce more verifiable facts. From the table 2, we observe a consistent decrease in NEIP for the TG-DRG-R compared to KAPING and BM25. Furthermore, we noticed a marked improvement in NEIP with TG-DRG-NR. This suggests that reformulating the dialogue can substantially enhance factuality.
Notably, the F1 score is not always consistent with the fact score. For instance, the trend of the F1 score shown in Flan-T5-XXL is almost opposite to the dialogue fact score on the OpendialKG dataset. The reasons can be attributed to either the reference responses lack explicit answers or the F1 score ignores the semantic aspects, and two sentences with the same entities can express contrary meanings.
In short, knowledge-augmented methods improve the factuality of LLMs, including the dialogue fact score and NEIP. Our proposed frameworks perform better than G-Retriever, KAPING and BM25 in the above metrics.
@@
With the KG, the BLEU-4 and ROUGE-L scores fluctuate across two datasets. Our proposed TG-DRG-NR generally perform better than other knowledge-enhanced baselines, while we see a drop with TG-DRG-R. Our proposed GA-DRG achieves the highest BLEU and ROUGE-L scores, surpassing the SOTA G-retriever and demonstrating the superiority of our method.
There is a slight growth in PPL of smaller language models like ChatGLM and Flan-T5-Large: incorporating the KG decreases text fluency. After analysis, we found that the relatively smaller LLMs sometimes can not understand the KG well and thus generate unpolished responses.
Overall, knowledge-augmented methods show variable performance across different LLMs in terms of BLEU-4 and ROUGE-L scores, and they tend to increase PPL slightly for smaller language models. Our proposed GA-DRG gains the best performance in BLEU-4 on both datasets and ROUGE-L on HybriDialogue, indicating a higher similarity between responses and references.
@@
We conduct experiments with different scales of LLMs. The Flan-T5 series LLMs are beneficial in analysing the correlation between size and performance.
As model scale grows, dialogue fact score and NEIP show differing trends: factuality declines on HybriDialogue but improves on OpendialKG, which contradicts our intuition that larger models ensure consistent factuality.
OpendialKG HybriDialogue
Methods $\downarrow Fact NEIP\downarrow$
w/o Know. Sel. -1.1 +1.2 -1.1 -0.11
w/o Dial. Reform. -1.46 +0.44 -2.3 +0.48
w/o Graph Repr. -1.00 -1.52 -3.41 -0.03
Table 3: An ablation study on GA-DRG, where “w/o” denotes the removal of components: Knowledge Selection (Know. Sel.), Dialogue Reformulation (Dial. Reform.), and Graph Representation (Graph Repr.).
@@ 4.7 Ablation Study
We conduct an ablation study with GA-DRG, shown in Table 3. We remove our components one by one from GA-DRG to observe the performance. We only report our aim—factually related metrics. The results show a noticeable decline in performance after removing each component, underscoring their critical roles in maintaining the factual accuracy of GA-DRG.
@@ 4.8 Human Evaluation Results
ChatGLM-6B 1.99 1.91 1.11 1.90 1.99 1.71
+BM25 1.99 1.89 1.09 1.85 1.97 1.75
+KAPING 1.98 1.95 1.18 1.84 1.98 1.71
+TG-DRG-NR 2.00 1.94 1.11 1.93 1.98 1.76
+TG-DRG-R 1.98 1.97 1.28 1.94 1.96 1.85
Flan-T5-XXL 1.56 2.00 1.36 1.66 1.69 1.74
+BM25 1.38 1.94 1.49 1.68 1.71 1.80
+KAPING 1.46 1.99 1.53 1.58 1.78 1.77
+TG-DRG-NR 1.52 1.98 1.39 1.62 1.76 1.77
+TG-DRG-R 1.47 1.97 1.65 1.55 1.85 1.76
G-retriever 1.96 1.96 0.71 1.92 1.82 1.84
GA-DRG 2.00 2.00 0.60 1.96 1.81 1.82
@@
We present the human evaluation results from three perspectives: coherence, fluency, and informativeness, as shown in Table 4. Each aspect is assessed on a three-level scale: positive (2 points), neutral (1 point), or negative (0 points) (e.g., coherent, neutral, incoherent). The average scores for each aspect are then calculated based on these ratings.
Our proposed frameworks TG-DRG and GA-DRG demonstrate marked improvements in coherence across various settings, with the exception of the Flan-T5-XXL model. In terms of fluency, our methods achieve the highest scores for Flan-T5-XXL on HybriDialogue, as well as for ChatGLM and graph-based methods on OpendialKG. Overall, the proposed methods maintain strong performance in fluency and coherence. Additionally, the TG-DRG generally enhance informativeness.
Case 1
Dialogue {Original / Coreference Resolved}:
Speaker A: What do you think about Wladimir Klitschko?
Speaker B: Not a whole lot but I do know {he / Wladimir Klitschko} is married to Hayden Panettiere from the show Nashville.
Speaker A: I have watched {that / Nashville} before, is {she / Hayden Panettiere} in any other show?
Speaker B: {She / Hayden Panettiere} played a voice on the cartoon A Bug’s Life. Other than that I can’t think of where I’ve seen {her / Hayden Panettiere}.
Speaker A: Ok, I didn’t know {she / Hayden Panettiere} voiced a character in {that / A Bug’s Life}.
Reference: I was surprised to hear that Kevin Spacey also voiced a character in that movie.
Flan-T5-XXL: She voiced a character in the cartoon A Bug’s Life.
BLEU-4: 13.04; ROUGE-L: 30.8
Key Knowledge: (Dot, voice actor, Hayden Panettiere)
TG-DRG-R: Hayden Panettiere voiced the character Dot in A Bug’s Life.
BLEU-4: 0; ROUGE-L: 23.1
Table 5: Dialogue and LLMs’ responses related to Hayden Panettiere, before and after coreference resolution.